Determine the amount of an immediate annuity of N$200 per annum left unpaid for 10 year at 7% p.a compounded interest.


1
Expert's answer
2021-05-07T11:02:01-0400

"FV=\frac{P[(1+r)^t-1]}{r}"

Now, substituting the values,

"\Rightarrow FV=\frac{200[(1+0.07)^10-1]}{0.07}"


"\Rightarrow FV= \frac{200(1.07^{10}-1)}{0.07}"


"\Rightarrow FV=2763.28"

Hence the required amount will be $2763.28